Minutes of nanoBPM meeting on 9 September, 2004

"Status of mover system", Y.Honda

(transparencies, 13 pages, pdf, 1.4MB )

Three sets of movers have been assembled and aligned in a series of #2, #1 and #3 on the granite table, where #2 is the repaired one. Performances of the movers have been measured. Hysteresis of all the movers was clearly observed with applied voltage from 0 to 140V. Dynamic moving range and cross talking are summarized in following tables for #2 and #3, respectively. Results of #1 can be seen at the previous meeting (2/19,2004).

Mover#2XX'YY'
unitμmμradμmμrad
X545.830.9
X'6300.9
Y20.5640.9
Y'33.853
Mover#3XX'YY'
unitμmμradμmμrad
X605.80.90.5
X'9400.9
Y10.5601.4
Y'04.8107

The resonance frequencies of assembled movers have been measured by hammering method. The typical damping time of oscillation was 100-200msec. The results are listed below;
setXYX'Y'
HzHzHzHz
182238139313
274222147313
378238164370
note: The results of the last meeting were those of individual movers.

Also, he explained a design of the V-block for the cavity-BPM base. The V-block is separated into two parts along beam axis to avoid a deformation of the sensor cavity. Surfaces of 1mm thickness were kept for future precise/final machining.

"Status of KEK cavity BPM", Y. Inoue and H.Hayano

The "beads" measurement of the first cavity-BPM was performed. Two resonant peaks were observed by network analyser. Their origin was not understood. The frequencies will be measured as a function of the beads position.

"Discussion on schedule"

We reminded the schedule in this month;

Actually, the de-installation of OTR was completed on 9 September. So, the floor can be clearly seen. Yamaoka took pictures as follows;

Next meeting will be held in 30 September, 2004, at 1:30pm-, 3-gokan 325.